# DepthSplat: Connecting Gaussian Splatting and Depth

### v1

@@ 3.4 Gaussian Parameter Prediction
For the task of 3D Gaussian splatting, we directly unproject the per-pixel depth maps to 3D with the camera parameters as the Gaussian centers $\bm{\mu}_{j}$. We append an additional lightweight network to predict other Gaussian parameters $\alpha_{j},\bm{\Sigma}_{j},\bm{c}_{j}$, which are opacity, covariance, and color, respectively. With all the predicted 3D Gaussians, we can render novel view images with the Gaussian splatting operation (Kerbl et al., 2023).
@@ 3.5 Training Loss
We study the properties of our proposed model on two tasks: depth estimation and novel view synthesis with 3D Gaussian splatting (Kerbl et al., 2023). The loss functions are introduced below.
@@
$L_{\mathrm{depth}}=\alpha\cdot|{\bm{D}}_{\mathrm{pred}}-{\bm{D}}_{\mathrm{gt}}|+\beta\cdot(|\partial_{x}{\bm{D}}_{\mathrm{pred}}-\partial_{x}{\bm{D}}_{\mathrm{gt}}|+|\partial_{y}{\bm{D}}_{\mathrm{pred}}-\partial_{y}{\bm{D}}_{\mathrm{gt}}|),$ (1)
where $\partial_{x}$ and $\partial_{y}$ denotes the gradients on the $x$ and $y$ directions, respectively. Following UniMatch (Xu et al., 2023), we use $\alpha=20$ and $\beta=20$.
View synthesis. We train our full model with a combination of mean squared error (MSE) and LPIPS (Zhang et al., 2018) losses between rendered and ground truth image colors:
@@
where $M$ is the number of novel views to render in a single forward pass. The LPIPS loss weight $\lambda$ is set to 0.05 following MVSplat (Chen et al., 2024).
@@ 4 Experiments
Implementation Details. We implement our method in PyTorch (Paszke et al., 2019) and optimize our model with the AdamW (Loshchilov & Hutter, 2017) optimizer and cosine learning rate learning rate schedule with warm up in the first $5\%$ of the total training iterations. We adopt the xFormers (Lefaudeux et al., 2022) library for our monocular ViT backbone implementation. We use a lower learning rate $2\times 10^{-6}$ for the pre-trained monocular backbone, and other remaining layers use a learning rate of $2\times 10^{-4}$. The feature downsampling factor $s$ in our multi-view branch (Sec. 3.1) is chosen based on the image resolution. More specifically, for experiments on the $256\times 256$ resolution RealEstate10K (Zhou et al., 2018) dataset, we choose $s=4$. For higher resolution datasets (e.g., TartanAir (Wang et al., 2020), ScanNet (Dai et al., 2017), KITTI (Geiger et al., 2013) and DL3DV (Ling et al., 2023)), we choose $s=8$. Our hierarchical matching models in Sec. 3.3 use 2-scale features, i.e., $1/8$ and $1/4$, or $1/4$ and $1/2$ resolutions.
Training Details. For depth experiments, we mainly follow the setup of UniMatch (Xu et al., 2023) for fair comparisons. More specifically, we train our depth model on 4 GH200 GPUs for 100K iterations with a total batch size of 32. For Gaussian splatting experiments on RealEstate10K (Zhou et al., 2018), we use the same training and testing splits of pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024), and train our model on 8 GH200 GPUs for 150K iterations with a total batch size of 32, which takes about 1 day. For experiments on the DL3DV (Ling et al., 2023) dataset, we evaluate on the official benchmark split with 140 scenes, while other remaining 9896 scenes are used for training. We fine-tune our RealEstate10K pre-trained model on 4 A100 GPUs for 100K iterations with a total batch size of 4, where the number of input views is randomly sampled from 2 to 6. We evaluate the model’s performance on different number of input views (2, 4, 6). Our code and training scripts are available at https://github.com/cvg/depthsplat.
Table 1: DepthSplat model variants. We explore different monocular backbones and different multi-view models ($1$-scale and 2-scale features for hierarchical matching as described in Sec. 3.3), where both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both depth estimation and view synthesis tasks.
Monocular Multi-View Depth (TartanAir) 3DGS (RealEstate10K) Param (M)
@@
ViT-S $1$-scale 8.46 93.02 26.76 0.877 0.123 37
ViT-B $1$-scale 6.94 94.46 27.09 0.881 0.119 113
ViT-L $1$-scale 6.07 95.52 27.34 0.885 0.118 354
ViT-S $2$-scale 7.01 94.56 26.96 0.880 0.122 40
ViT-B $2$-scale 6.22 95.31 27.27 0.885 0.120 117
ViT-L $2$-scale 5.57 96.07 27.44 0.887 0.119 360
Table 2: Ablations. We evaluate the contribution of the monocular feature branch and the cost volume branch, as well as different monocular features. Our results indicate that the monocular feature and cost volume are complementary, with large performance drops when removing either one. The pre-trained Depth Anything network weights achieve overall the best performance.
Module Method Depth (TartanAir) 3DGS (RealEstate10K)
@@ 4.1 Model Variants
We first study several different model variants for both depth estimation and view synthesis tasks. In particular, we explore different monocular backbones (Yang et al., 2024b) (ViT-S, ViT-B, ViT-L) and different multi-view models ($1$-scale and $2$-scale). We conduct depth experiments on the large-scale TartanAir (Wang et al., 2020) synthetic dataset, which features both indoor and outdoor scenes and has perfect ground truth depth. The Gaussian splatting experiments are conducted on the standard RealEstate10K (Zhou et al., 2018) dataset. Following community standards, we report the depth evaluation metrics (Eigen et al., 2014) of Abs Rel (relative $\ell_{1}$ error) and $\delta_{1}$ (percentage of correctly estimated pixels within a threshold) and novel view synthesis metrics (Kerbl et al., 2023) of PSNR, SSIM, and LPIPS. The results in Table 1 demonstrate that both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both tasks.
@@
From Table 1, we can also see that better depth network architecture leads to improved view synthesis. In Appendix A.1, we conduct additional experiments to study the effect of different initializations for the depth network to the view synthesis performance. Our results show that a better depth model initialization also contributes to improved rendering quality. In summary, both better depth network architecture and better depth model initialization lead to improved novel view synthesis results.
4.2 Ablation Study and Analysis
@@
In this section, we study the properties of our key components on the TartanAir dataset (for depth estimation) and RealEstate10K dataset (for view synthesis with Gaussian splatting).
@@
Monocular Features for Depth Estimation and Gaussian Splatting. In Table 2, we compare our full model (full) with the model variant where the monocular depth feature branch (with a ViT-S model pre-trained by Depth Anything v2 (Yang et al., 2024b)) is removed (w/o mono feature), leaving only the multi-view branch. We can observe a clear performance drop for both depth and view synthesis tasks. In Fig. 3, we visualize the depth predictions and error maps of both models on the ScanNet dataset. The pure multi-view feature matching-based approach struggles a lot at texture-less regions and reflective surfaces. At the same time, our full model achieves reliable results thanks to the powerful prior captured in monocular depth features. We also show the visual comparisons for the Gaussian splatting task in Fig. 4 with two input views. For regions (e.g., the lounge chairs) that only appear in a single image, the pure multi-view method is unable to find correspondences. It thus produces unreliable depth predictions, leading to misalignment in the rendered novel views due to the incorrect geometry.
@@
We also experiment with removing the cost volume (w/o cost volume) in the multi-view branch and observed a significant performance drop. This indicates that obtaining scale- and multi-view consistent predictions with a pure monocular depth backbone is challenging, which constrains achieving high-quality results in downstream tasks.
@@
Fusion Strategy. We compare with alternative strategies for fusing the monocular features for multi-view depth estimation. In particular, we compare with MVSFormer (Cao et al., 2022) which constructs the cost volume with monocular features. More specifically, we replace our multi-view feature extractor with a weight-sharing ViT model and use the ViT features to build the cost volume as done in MVSFormer. This leads to a single-branch architecture (single branch in Table 2), unlike our two-branch design where the monocular features are not used to build the cost volume. We can observe that our fusion strategy performs significantly better than the single-branch design, potentially because our two-branch design disentangles feature matching and obtaining monocular priors, which makes the learning task easier.
Different Monocular Features. In Table 2, we also evaluate different monocular features, including the ConvNeXt-T (Liu et al., 2022) features used in AFNet (Cheng et al., 2024), and other popular monocular features including Midas (Ranftl et al., 2020) and DINO v2 (Oquab et al., 2023). The pre-trained Depth Anything monocular features achieve overall the best performance.
Table 3: Unsupervised Depth Pre-Training with Gaussian Splatting. We first perform unsupervised pre-training with Gaussian splatting on RealEstate10K and then measure the depth performance on TartanAir, ScanNet and KITTI after fine-tuning for the depth task. Compared to previous supervised pre-trained models Depth Anything (for monocular ViT) and UniMatch (for multi-view Transformer), our new unsupervised pre-training improves performance consistently in all metrics. The improvements are especially significant on the challenging datasets like TartanAir and KITTI.
Initialization TartanAir ScanNet KITTI
@@
Depth Anything (only mono) 11.12 89.97 6.78 96.13 11.67 85.96
UniMatch + Depth Anything (mv & mono) 10.86 90.55 6.70 96.14 11.56 87.27
DepthSplat (full model) 10.20 91.10 6.60 96.27 10.68 89.92
4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
By connecting Gaussian splatting and depth, our DepthSplat provides a way to pre-train the depth model in a fully unsupervised manner. In particular, we first train our full model on the large-scale unlabelled RealEstate10K dataset (contains $\sim$ 67K Youtube videos) with only the Gaussian splatting rendering loss (Eqn. 2), without any direct supervision on the depth predictions. After pre-training, we take the pre-trained depth model and further fine-tune it to the depth task on the mixed TartanAir and VKITTI2 (Cabon et al., 2020) datasets with ground truth depth supervision.
In Table 3, we evaluate the performance on both in-domain TartanAir test set and the zero-shot generalization on unseen ScanNet and KITTI datasets. We compare with only initializing the monocular backbone with Depth Anything and additionally initializing the multi-view Transformer backbone with UniMatch (Xu et al., 2023). Our approach achieves the best results on all three datasets. It is also interesting to observe that our pre-training contributes more on more challenging datasets (i.e., TartanAir and KITTI which feature complex large-scale scenes unlike ScanNet only contains indoor scenes). We also note that both Depth Anything and UniMatch are trained with ground truth supervision, while our DepthSplat is trained with only photometric losses. Given the increasing popularity of view synthesis (Zheng & Vedaldi, 2024; Weng et al., 2023) and multi-view generative models (Shi et al., 2023; Blattmann et al., 2023), new multi-view datasets (Ling et al., 2023) and models (Voleti et al., 2024) are gradually introduced, our approach provides a way to pre-train depth models on large-scale unlabelled multi-view image datasets. This could potentially further improve the multi-view consistency and robustness of existing depth models (Yang et al., 2024a; Piccinelli et al., 2024), which are usually trained with ground truth depth supervision.
4.4 Benchmark Comparisons
Comparisons on ScanNet and RealEstate10K. Table 5 and Table 5 compare the depth and novel view synthesis results on the standard ScanNet and RealEstate10K benchmarks, respectively. For both comparisons, the numbers of input images are two. We can observe that our DepthSplat outperforms previous methods Teed & Deng (2019); Yu et al. (2021); Xu et al. (2024a); Chen et al. (2024) by clear margins on both datasets for both tasks. The visual comparison with previous best methods is shown in Fig. 5 and Fig. 6, where our method significantly improves the performance on challenging scenarios like texture-less regions and occlusions.
Comparisons on DL3DV. To further evaluate the performance on complex real-world scenes, we conduct comparisons with the latest state-of-the-art Gaussian splatting model MVSplat (Chen et al., 2024) on the recently introduced DL3DV dataset (Ling et al., 2023). We also compare the results of different numbers of input views (2, 4 and 6) on this dataset. We fine-tune MVSplat and our RealEstate10K pre-trained models on DL3DV training scenes and report the results on the benchmark test set in Table 6. Our DepthSplat consistently outperforms MVSplat in all metrics. Visual comparisons with MVSplat on the DL3DV dataset are shown in Fig. 7, where MVSplat’s depth quality lags behind our DepthSplat due to matching failure, which leads to blurry and distorted view synthesis results. We show more visual comparison results in Appendix A.2. It is also worth noting that our method scales more efficiently to more input views thanks to our lightweight local feature matching approach (Sec 3.1), which is unlike the expensive pair-wise matching used in MVSplat.
We invite the readers to our project page https://haofeixu.github.io/depthsplat/ for the video results on different number of input views (6 and 12), where our DepthSplat is able to reconstruct larger-scale or 360 scenes from more input views, without any optimization.
@@ 4.4 Benchmark Comparisons
DeMoN 0.231 0.761 0.289
BA-Net 0.161 0.346 0.214
DeepV2D 0.057 0.168 0.080
NeuralRecon 0.047 0.164 0.093
DRO 0.053 0.168 0.081
UniMatch 0.059 0.179 0.082
DepthSplat (ours) 0.044 0.119 0.059
@@
pixelNeRF 20.43 0.589 0.550
GPNR 24.11 0.793 0.255
AttnRend 24.78 0.820 0.213
MuRF 26.10 0.858 0.143
pixelSplat 25.89 0.858 0.142
MVSplat 26.39 0.869 0.128
DepthSplat (ours) 27.44 0.887 0.119
Figure 5: Depth Comparison on ScanNet. Our DepthSplat performs significantly better than UniMatch (Xu et al., 2023) on challenging parts (e.g., edges of the couch and pillows).
Figure 6: Visual Synthesis on RealEstate10K. Our DepthSplat performs significantly better than pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024) in challenging regions.

### v2

@@ 3.4 Gaussian Parameter Prediction
For the task of 3D Gaussian splatting, we directly unproject the per-pixel depth maps to 3D with the camera parameters as the Gaussian centers $\bm{\mu}_{j}$. To predict other remaining Gaussian parameters $\alpha_{j}$ (opacity), $\bm{\Sigma}_{j}$ (covariance) and $\bm{c}_{j}$ (color), we use an additional lightweight 2D U-Net [39], where it takes the concatenated image, depth and feature information as input and outputs all the remaining Gaussian parameters. With all the predicted 3D Gaussians, we can render novel views with the Gaussian splatting operation [27].
@@ 3.5 Training Loss
We study the properties of our proposed model on two tasks: depth estimation and novel view synthesis with 3D Gaussian splatting [27]. The loss functions are introduced below.
@@
$L_{\mathrm{depth}}$ $=\alpha\cdot|{\bm{D}}_{\mathrm{pred}}-{\bm{D}}_{\mathrm{gt}}|+\beta\cdot|\partial_{x}{\bm{D}}_{\mathrm{pred}}-\partial_{x}{\bm{D}}_{\mathrm{gt}}|$
$\quad+\beta\cdot|\partial_{y}{\bm{D}}_{\mathrm{pred}}-\partial_{y}{\bm{D}}_{\mathrm{gt}}|),$ (1)
where $\partial_{x}$ and $\partial_{y}$ denotes the gradients on the $x$ and $y$ directions, respectively. Following UniMatch [61], we use $\alpha=20$ and $\beta=20$.
View synthesis. We train our full model with a combination of mean squared error (MSE) and LPIPS [73] losses between rendered and ground truth image colors:
@@
where $M$ is the number of novel views to render in a single forward pass. The LPIPS loss weight $\lambda$ is set to 0.05 [8].
@@ 4 Experiments
Implementation Details. We implement our method in PyTorch [35] and optimize our model with the AdamW [33] optimizer and cosine learning rate schedule. We adopt the xFormers [28] library for our monocular ViT backbone implementation. We use a lower learning rate $2\times 10^{-6}$ for the pre-trained Depth Anything V2 [66] backbone, and other remaining layers use a learning rate of $2\times 10^{-4}$. The feature downsampling factor $s$ in our multi-view branch (Sec. 3.1) is chosen based on the image resolution. More specifically, for experiments on the $256\times 256$ resolution RealEstate10K [75] dataset, we choose $s=4$. For higher resolution datasets (e.g., TartanAir [55], ScanNet [12], KITTI [22] and DL3DV [30]), we choose $s=8$. Our hierarchical matching models in Sec. 3.3 use 2-scale features, i.e., $1/8$ and $1/4$, or $1/4$ and $1/2$ resolutions.
Training Details. For depth experiments, we mainly follow the setup of UniMatch [61] for fair comparisons on the ScanNet [12] dataset. More specifically, we train our depth model on 4 GH200 GPUs for 100K iterations with a total batch size of 32. For the ablation experiments on depth task, we mainly use synthetic datasets (TartanAir [55] and VKITTI2 [5]) for training since they provide high-quality ground truth depths. For Gaussian splatting experiments, we consider both low-resolution and high-resolution images. In particular, for comparisons with previous methods [7, 8], we mainly use $256\times 256$ resolution RealEstate10K [75] and $256\times 448$ resolution DL3DV [30] datasets. We also report high-resolution ($512\times 960$) results on RealEstate10K and DL3DV datasets for qualitative evaluations. For experiments on the $256\times 256$ RealEstate10K dataset, we train our model on 8 GH200 GPUs for 150K iterations with a total batch size of 32, which takes about 20 hours. For experiments on the $256\times 448$ DL3DV [30] dataset, we evaluate on the official benchmark split with 140 scenes, while other remaining scenes are used for training. We fine-tune our RealEstate10K pre-trained model on 4 A100 GPUs for 100K iterations with a total batch size of 4, where the number of input views is randomly sampled from 2 to 6. We evaluate the model’s performance on different number of input views (2, 4, 6). For high-resolution results, we fine-tune our low-resolution pre-trained models on high-resolution images. More training details are provided in the appendix. Our code and training scripts are available at github.com/cvg/depthsplat to ease reproducibility.
Table 1: DepthSplat model variants. We evaluate different monocular backbones and different multi-view models ($1$-scale and 2-scale features for hierarchical matching as described in Sec. 3.3), where both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both depth and view synthesis tasks. Our model parameters, inference time, and memory consumption are provided in Tab. A.1 of the appendix.
Monocular Multi-View Depth (TartanAir) 3DGS (RealEstate10K)
@@
ViT-S $1$-scale 8.46 93.02 26.76 0.877 0.123
ViT-B $1$-scale 6.94 94.46 27.09 0.881 0.119
ViT-L $1$-scale 6.07 95.52 27.34 0.885 0.118
ViT-S $2$-scale 7.01 94.56 26.96 0.880 0.122
ViT-B $2$-scale 6.22 95.31 27.27 0.885 0.120
ViT-L $2$-scale 5.57 96.07 27.44 0.887 0.119
Table 2: Ablations. We evaluate the contribution of the monocular feature branch and the cost volume branch (1st group of experiments), as well as different monocular features (2nd group of experiments). Our results indicate that the monocular feature and cost volume are complementary, with large performance drops when removing either one. The pre-trained Depth Anything V2 model weights achieve the best view synthesis results.
Components Depth (TartanAir) 3DGS (RealEstate10K)
@@ 4.1 Model Variants
We first study several different model variants for both depth estimation and view synthesis tasks. In particular, we explore different monocular backbones [66] (ViT-S, ViT-B, ViT-L) and different multi-view models ($1$-scale and $2$-scale). We conduct depth experiments on the large-scale TartanAir [55] synthetic dataset, which features both indoor and outdoor scenes and has perfect ground truth depth. The Gaussian splatting experiments are conducted on the standard RealEstate10K [75] dataset. Following community standards, we report the depth evaluation metrics [18] of Abs Rel (relative $\ell_{1}$ error) and $\delta_{1}$ (percentage of correctly estimated pixels within a threshold) and novel view synthesis metrics [27] of PSNR, SSIM, and LPIPS. The results in Tab. 1 demonstrate that both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both tasks.
@@
From Tab. 1, we can also observe that better depth network architecture leads to improved view synthesis. In Tab. B.1 of the appendix, we conduct additional experiments to study the effect of different initializations for the depth network to the view synthesis performance. Our results show that a better depth model initialization also contributes to improved rendering quality. In summary, both better depth network architecture and better depth model initialization lead to improved novel view synthesis results.
4.2 Ablation and Analysis
@@
In this section, we study the properties of our key components on the TartanAir dataset (for depth) and RealEstate10K dataset (for view synthesis with Gaussian splatting).
@@
Monocular Features for Depth and Gaussian Splatting. In Tab. 2, we compare our full model (full) with the model variant where the monocular depth feature branch (with a ViT-S model pre-trained by Depth Anything V2 [66]) is removed (w/o mono feature), leaving only the multi-view branch. We can observe a clear performance drop for both depth and view synthesis tasks. In Fig. 3, we visualize the depth predictions and error maps of both models on the ScanNet dataset. The pure multi-view feature matching-based approach struggles a lot at texture-less regions and reflective surfaces. At the same time, our full model achieves reliable results thanks to the powerful prior captured in monocular depth features. We also show the visual comparisons for the Gaussian splatting task in Fig. 4 with two input views. For regions (e.g., the lounge chairs) that only appear in a single image, the pure multi-view method is unable to find correspondences. It thus produces unreliable depth predictions, leading to misalignment in the rendered novel views due to the incorrect geometry.
@@
We also experiment with removing the cost volume (w/o cost volume) in the multi-view branch and observed a significant performance drop. This indicates that obtaining scale- and multi-view consistent predictions with a pure monocular depth backbone is challenging, which constrains achieving high-quality results for 3D Gaussian reconstruction.
@@
Fusion Strategy. We compare with alternative strategies for fusing the monocular features for multi-view depth estimation. In particular, we compare with MVSFormer [6] which constructs the cost volume with monocular features. More specifically, we replace our multi-view feature extractor with a weight-sharing ViT model and use the ViT features to build the cost volume as done in MVSFormer. This leads to a single-branch architecture (single branch in Tab. 2), unlike our two-branch design where the monocular features are not used to build the cost volume. We can observe that our fusion strategy performs significantly better than the single-branch design, potentially because our two-branch design disentangles feature matching and obtaining monocular priors, which makes the learning task easier.
Table 3: Unsupervised Depth Pre-Training with Gaussian Splatting. We evaluate the same architecture but with different weights. The pre-training is performed on RealEstate10K, and fine-tuning is performed on TartanAir and VKITTI2. Unsupervised pre-training (2nd row) produces improved depth predictions compared to random initialization (1st). The performance can be further improved with additional fine-tuning (4th), where the benefit of pre-training is especially significant on challenging datasets like TartanAir and KITTI compared to training from random initialization (3rd).
Pre-Train Fine-Tune TartanAir ScanNet KITTI
@@
76.22 3.43 39.97 22.56 90.69 0.00
29.53 53.16 21.51 76.09 56.83 46.26
10.86 90.55 6.70 96.14 11.56 87.27
10.20 91.10 6.60 96.27 10.68 89.92
@@
DeMoN [51] 0.231 0.761 0.289
BA-Net [47] 0.161 0.346 0.214
DeepV2D [49] 0.057 0.168 0.080
NeuralRecon [44] 0.047 0.164 0.093
DRO [24] 0.053 0.168 0.081
UniMatch [61] 0.059 0.179 0.082
DepthSplat 0.044 0.119 0.059
Different Monocular Features. In Tab. 2, we also evaluate different monocular features, including the ConvNeXt-T [32] features used in AFNet [9], and other popular monocular features including Midas [37] and DINOv2 [34]. The pre-trained Depth Anything V2 [66] monocular features achieve the best view synthesis results.
4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
By connecting Gaussian splatting and depth, our DepthSplat provides a way to pre-train the depth model in a fully unsupervised manner. In particular, we first train our full model on the large-scale unlabeled RealEstate10K dataset (contains $\sim$ 67K Youtube videos) with only the Gaussian splatting rendering loss (Eq. 2), without any direct supervision on the depth predictions. After pre-training, we take the pre-trained depth model and further fine-tune it to the depth task on the mixed TartanAir [55] and VKITTI2 [5] datasets with ground truth depth supervision. In Tab. 3, we evaluate the performance on both in-domain TartanAir test set and the zero-shot generalization on unseen ScanNet and KITTI datasets. We can observe that the benefit of pre-training is especially significant on the challenging datasets like TartanAir and KITTI. The results align well with the observation in [19] which suggests that the pre-training acts as regularization and guides the learning towards better minima and enables better generalization to out-of-distribution datasets. The visual comparison results are provided in Fig. C.1 of the appendix. Given the increasing popularity of view synthesis [74, 57] and multi-view generative models [42, 3], new multi-view datasets [30] and models [52] are gradually introduced, our approach provides a way to pre-train depth models on large-scale unlabeled multi-view image datasets.
@@ 4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
pixelNeRF [70] 20.43 0.589 0.550
GPNR [43] 24.11 0.793 0.255
AttnRend [15] 24.78 0.820 0.213
MuRF [62] 26.10 0.858 0.143
pixelSplat [7] 25.89 0.858 0.142
MVSplat [8] 26.39 0.869 0.128
TranSplat [71] 26.69 0.875 0.125
DepthSplat 27.44 0.887 0.119
Figure 5: Visual synthesis on RealEstate10K. Our DepthSplat performs significantly better than pixelSplat [7] and MVSplat [8] in challenging regions.
Figure 6: Depth results on unseen samples. Our depth model generalizes robustly to different scene types.
4.4 Benchmark Comparisons
Comparisons on ScanNet and RealEstate10K. Tab. 4 and Tab. 5 compare the depth and novel view synthesis results on the standard ScanNet and RealEstate10K benchmarks, respectively. We can see clearly that our DepthSplat achieves state-of-the-art performance on both datasets for both tasks. The visual comparison with previous methods is shown in Fig. 5, where our method significantly improves the performance on challenging scenarios like texture-less regions and occlusions.